40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 5 21:42:28 2009, Yoichi, Update, PSL, My thoughts on ISS INMONPD_Spectrum_1-10Hz.pdf
    Reply  Fri Feb 6 09:52:35 2009, Kakeru, Update, PSL, Current shunt transfar function TF_CS_gain.pngTF_CS_phase.png
       Reply  Fri Feb 6 23:23:48 2009, Kakeru, Yoichi, Update, PSL, ISS is fixed Current_ShuntTF_gain.pngCurrent_ShuntTF_phase.png
          Reply  Mon Feb 9 16:02:42 2009, Yoichi, Update, PSL, PSL relative intensity noise RIN-13dB.pngRIN-21dB.png
             Reply  Tue Feb 10 23:36:25 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
                Reply  Wed Feb 11 07:28:25 2009, Yoichi, Update, PSL, PA current and laser output 
                   Reply  Wed Feb 11 23:51:53 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
Message ID: 1277     Entry time: Fri Feb 6 09:52:35 2009     In reply to: 1276     Reply to this: 1283
Author: Kakeru 
Type: Update 
Category: PSL 
Subject: Current shunt transfar function 

I attach the transfar function of the current shunt.
There is a little gap at 10 Hz for phase, but it is a ploblem of measurement and not real one.
 

Attachment 1: TF_CS_gain.png  12 kB  | Show | Show all
Attachment 2: TF_CS_phase.png  10 kB  | Show | Show all
ELOG V3.1.3-